At Ed Gas, regular gas is $1.12 per gallon and ultra gas is $1.84 per gallon. If Ed Gas sold $128.80 worth of ultra, how many gallons of regular must be sold to have an average selling price of $1.518 per gallon?
Answer by mananth(16946)

At Ed Gas, regular gas is $1.12 per gallon
ultra gas is $1.84 per gallon. $128.80 worth of ultra was sold,
Let thenumber of gallons of regular to be sold be x to have an average selling price of $1.518 per gallon?
..
Ultra 128.80 : price = 1.84/gallon
Number of gallons of ultra = 128.80/1.84 = 70 gallons
..
Value of regular + value of ultra / number of gallons = average price
1.12x+128.80 /70+x= 1.518
1.12x+128.80 =(70+x)* 1.518
1.12x+128.80=1.518x+106.26
1.518x-1.12x=106.26-128.80
0.398x=22.54
=56.63 gallons of regular has to be sold